Compute the discount date, the due date, the amount of discount and amount of the remittance if the required remittance is paid within the discount period. Round your answers to two decimal places.

Terms: 2/10, net/25 Discount date: Select: Oct. 7, Oct. 23, Nov. 7, Dec. 23
Invoice date: Oct. 13 Due date: Select: Oct. 7, Oct. 23, Nov. 7 Dec. 23
Invoice amount: $819.24 Discount amount: $
Freight charge: $52.12 Remittance: $
Answer #1

Due date = Oct 23

Discount amount = 819.25 - 2% = $16.39

Remittance = 819.25 + 52.12 - 16.39 = $854.98